OpenStudy (anonymous):

If g(x) = 2(x - 4), find the value of x if g(x) = 20. 32 12 14 10

OpenStudy (anonymous):

32

OpenStudy (anonymous):

@Xmoses1 @djcool31

OpenStudy (anonymous):

yeah baby :)

OpenStudy (xmoses1):

Wait no it is 14. Substitute 14 for the x and solve it

OpenStudy (anonymous):

nuh uh y ?

OpenStudy (anonymous):

i thought it was 2(20-4)

OpenStudy (anonymous):

me 2

OpenStudy (xmoses1):

If g(x) = 20, then what will x have to equal if 2(x - 4)= 20

OpenStudy (anonymous):

2(20-4) 20-4=16 2(16) 16x16=32

OpenStudy (xmoses1):

You need the answer to come out as 20... Or wait am i missing a step @djcool31??

OpenStudy (xmoses1):

Always do what is in the parenthesis first

OpenStudy (anonymous):

yo i think 32 is the answer thats all i got bro

OpenStudy (anonymous):

i'll be back later tho i gtg

OpenStudy (xmoses1):

Wow i suck at explaining... @MathLegend could you check this??

OpenStudy (xmoses1):

Better yet @whpalmer4

OpenStudy (math&ing001):

g(x)=2(x-4) Find x when g(x)=20 We need to solve this equation : 20=2(x-4) 10=x-4 x=14 Have a nice day !

OpenStudy (xmoses1):

Aha i was right thanks ^^